find three consecutive even number if thier sum 966
Found 2 solutions by mananth, Alan3354:
Answer by mananth(16949)

let the numbers be x,x+2,x+4
x+x+2+x+4 =966
3x+6 = 966
3x= 966-6
3x=960
x=320
320,322,324 are the numbers

Answer by Alan3354(69443)

find three consecutive even number if thier sum 966
----------
966/3 = 322, the middle number.
---
Or
n-2 + n + n+2 = 966
3n = 966
etc